OpenStudy (anonymous):

Washington High School has a total of 90 boys and girls who play sports. If the number of boys is 10 more than three times the number of girls, how many girls play sports at this high school? Answer 30 20 70 80

OpenStudy (anonymous):

boy = b girl = g b + g = 90 or b = 90 - g b = 3g + 10 Substitute and solve. :)

OpenStudy (anonymous):

20?